OpenStudy (anonymous):

the point-slope form of the equation that represents the line that passes through the points (-6, 4) and (2, 0).

OpenStudy (anonymous):

(x1,y1) = (-6,4) (x2, y2) = (2,0) Point slope formula \[(y-y_1) = m(x-x_1)\] \[m = (y_2-y_1)/(x_2-x_1)\]

OpenStudy (anonymous):

Thanks so much!

OpenStudy (anonymous):

Anytime:)